A student is rotating a bucket of water in a vertical circle with a radius of 0.85m. In order for water to not spill out at the top of the swing, the bucket requires a minimum speed of ___m/s. (Give your answer with the correct number of significant digits and do no include units)


1
Expert's answer
2021-12-02T10:03:54-0500
"F_c=F_g,""\dfrac{mv^2}{r}=mg,""v=\sqrt{gr}=\sqrt{9.8\ \dfrac{m}{s^2}\times0.85\ m}=2.9\ \dfrac{m}{s}."